OpenStudy (anonymous):

TRIG PROBLEM: Check all that apply. If csc⊖ = 13/12, then: A) sec⊖= 12/13 B) sin⊖= 12/13 C) tan⊖=12/5 D) cos⊖= 12/13

geerky42 (geerky42):

\(\csc\theta = \dfrac{1}{\sin\theta}=\dfrac{Hyp}{Opp} = \dfrac{13}{12}\) So we can see that Hyp = 13 and Opp = 12: Now draw triangle: Now do you know how to find Adj?

OpenStudy (anonymous):

I'm lookin' through my study for this quiz and I haven't found anything that finds the adj... -_-

geerky42 (geerky42):

it's because the way to find Adj is what you learned in Algebra *cough* \(a^2+b^2=c^2\) *cough*

OpenStudy (anonymous):

and so the Pythagorean theorem returns.... I totally forgot about that